OpenStudy (anonymous):

What is the value of y in the equation 2(3y + 7 + 5) = 196 − 16?

OpenStudy (anonymous):

2(3y+7+5)=180 6y+24=180 6y=180-24 6y=156 y=156/6 y=???

OpenStudy (anonymous):

you can understand that?

OpenStudy (anonymous):

Yes thankyou

OpenStudy (anonymous):

yw